### 28 May 2026 v2

# Intersubjectivity as a principle determining physical observables and non-classicality

Abstract
We identify an operational principle that singles out Projection-Valued Measures (PVMs) among general Positive Operator-Valued Measures (POVMs), bridging the modern quantum measurement theory and the traditional formulation based on projective measurements of physical observables. We reformulate Ozawa’s intersubjectivity condition, which requires inter-observer agreement of the measurement outcomes, in a quantitative manner within the framework of generalized probabilistic theories. We prove that (i) a POVM is a PVM if and only if its every coarse-graining is intersubjective, and (ii) a system is classical if and only if intersubjectivity is preserved under any coarse-graining, establishing a complete characterization of the physical observables and the classical theory. Furthermore, measurements with intersubjectivity are sufficiently rich for the informational tasks of state tomography and state discrimination, testifying to its operational significance in quantum and beyond information processing.
Introduction.—
The discovery of quantum theory revealed the inherently probabilistic nature of the physical world, opening the way for measurement theory and an operational understanding of physical theories. Traditionally, quantum theory has been formulated in terms of physical observables represented by self-adjoint operators. In this framework, the measurement of an observable is described by a projective measurement obeying the Born rule. In contrast, modern quantum measurement theory adopts an operational perspective centered on Positive Operator-Valued Measures (POVMs), which encompass the most general class of physically allowed measurements. While many areas of physics still rely on the traditional approach, the operational framework has grown increasingly important, particularly with the rapid development of quantum information science.
From the modern perspective, the measurement of a physical observable is regarded as a special case of POVMs, namely, Projection-Valued Measures (PVMs). However, PVMs are defined in an algebraic way, and their operational characterization remains unclear. The aim of this Letter is to provide an operational characterization of PVMs among POVMs. In order to verify a genuinely operational interpretation of the results, we employ the framework of Generalized Probabilistic Theories (GPTs) [1, 2, 3, 4, 5, 6, 7, 8, 9], which retain only the operational requirements regarding states and measurements from the axioms of quantum theory while discarding the rest. Within GPTs, the generalization of projective measurements, and therefore also that of “physical observables,” is not yet clear, and this question remains an active topic of research [10, 11, 12, 13, 14, 15, 16, 17]. Numerous studies have also attempted to quantify the “PVM-likeness” of a POVM [18, 19, 20, 21, 22]; however, these approaches often lack a clear operational interpretation or fail to provide a complete characterization of PVMs.
In this Letter, we provide a fully operational characterization of PVMs along with their quantification by generalizing the concept of intersubjectivity [23, 24, 25] introduced by Ozawa to GPTs. A quantum measurement is called intersubjective if two observers performing the measurement simultaneously are guaranteed to obtain the same outcome. Ozawa argued that such inter-observer agreement is essential for a measurement to ascertain an observer-independent value, rather than merely generating outcomes randomly (Fig. 1). While Ozawa proved that every PVM is intersubjective in quantum theory [23], the converse implication remains an open question.
33
(a) Intersubjective measurement
14
(b) Randomly generating measurement
Figure 1: Conceptual diagram of intersubjectivity. Two observers simultaneously perform the same measurement. (a) An intersubjective measurement guarantees identical outcomes for both observers, whereas (b) a generic measurement may produce random outcomes independently.
We first show that intersubjectivity alone is insufficient to characterize PVMs. By comparing the intersubjectivity criterion with Gudder’s notion of sharp effects in GPTs [10, 11, 12] and specializing it to quantum theory, we derive a necessary and sufficient condition for a POVM to be intersubjective. We then construct explicit examples of intersubjective POVMs that are not PVMs. This resolves the converse direction left open in Ozawa’s analysis [23] and reveals a genuine gap between intersubjectivity and projectivity in quantum theory.
To fill the gap between intersubjective measurements and PVMs, we introduce the notion of a completely intersubjective measurement, namely, a measurement for which all of its coarse-grainings are intersubjective. We show that complete intersubjectivity provides a complete characterization of PVMs in quantum theory. Accordingly, complete intersubjectivity emerges as a natural candidate for the definition of “physical observables” in GPTs. Furthermore, complete intersubjectivity yields a quantification of “PVM-likeness” that carries a direct operational interpretation.
The gap between intersubjectivity and complete intersubjectivity indicates a paradoxical situation in which the objectivity of measurement outcomes can be lost when the resolution of the measurement is reduced. Remarkably, this counterintuitive phenomenon is common to all non-classical theories. We prove it rigorously by constructing measurements that are intersubjective but not completely intersubjective in any non-classical theory. Along the lines of research on GPTs that has presented characterizations of classical [26, 27, 28, 29, 30, 31, 32] and quantum [15, 33, 34, 35, 36, 37] theories, this result establishes a novel operational characterization of classical theory.
Finally, to further substantiate the operational significance of (complete) intersubjectivity, we show that (completely) intersubjective measurements are sufficiently rich in any GPT in terms of standard information-processing tasks, namely, state tomography [38, 39, 40] and state discrimination tasks [41, 42, 43, 37]. This demonstrates that (complete) intersubjectivity not only bridges the traditional observable-based formulation of quantum theory with the modern operational framework but also carries practical relevance for quantum information processing.
Generalized probabilistic theories.—
We briefly introduce generalized probabilistic theories (GPTs); for further details, see, e.g., [6, 7, 8, 9]. GPTs provide a framework based solely on operational axioms concerning states and measurements. This framework encompasses classical and quantum theories, as well as all physically conceivable theories beyond them [44, 45, 46, 47, 48, 49, 50].
In a GPT, a physical system is specified by its states and measurements. A state is represented by an element of a compact convex set $S$, called the state space, where the convex structure encodes probabilistic mixtures. A extreme point of $S$ is called a pure state, and we denote the set of all pure states by $\mathrm{ext}(S)$.
An affine map $a:S\to[0,1]$ is called an effect, and the set of all effects is denoted by $\mathcal{E}_{S}$. The state space is assumed to be separated by effects: for each $s_{1}\neq s_{2}\in S$, there exists an effect $a\in\mathcal{E}_{S}$ such that $a(s_{1})\neq a(s_{2})$. The partial order on $\mathcal{E}_{S}$ is defined by $a\leq b:\Leftrightarrow(a(s)\leq b(s)\text{ for all }s\in S)$. The constant maps $0_{S}$ and $1_{S}$, which return 0 and $1$ for all states, are the least and greatest elements of $\mathcal{E}_{S}$, respectively.
A finite-outcome measurement is given by a family $A=(a_{x})_{x\in X}$ of effects satisfying $\sum_{x\in X}a_{x}=1_{S}$. When a measurement $A=(a_{x})_{x\in X}$ is performed on a state $s\in S$, the probability of obtaining the outcome $x\in X$ is $a_{x}(s)$. We denote by $\mathcal{M}_{S}(X)$ the set of all measurements with outcome set $X$. We extend several of the results for continuous-outcome measurements; see the Supplemental Material  [^1] See Supplemental Material for further details. .
As standard examples, the state space of a classical system is given by a simplex, while that of a quantum system is given by the space of density operators on a Hilbert space. In quantum theory, effects correspond to positive operators bounded above by the identity operator, and measurements correspond to POVMs.
A measurement $C=(c_{x,y})_{(x,y)\in X\times Y}$ is called a joint measurement of a pair of measurements $A=(a_{x})_{x\in X}$ and $B=(b_{y})_{y\in Y}$ if its marginals reproduce $A$ and $B$, that is, $\sum_{y\in Y}c_{x,y}=a_{x}$ for all $x\in X$ and $\sum_{x\in X}c_{x,y}=b_{y}$ for all $y\in Y$. We denote the set of all joint measurements of $A$ and $B$ by $\mathrm{JM}(A,B)$. A pair of measurements $(A,B)$ is said to be compatible [52] if $\mathrm{JM}(A,B)\neq\emptyset$.
Intersubjectivity and sharpness.—
The concept of intersubjectivity plays a central role in this Letter. Ozawa [23] introduced intersubjectivity as a property of quantum observables requiring that all observers obtain the same outcome when they simultaneously measure the same observable (Fig. 1). The term “intersubjective” indicates that the measurement outcome is shared among observers, even though it is not assumed to represent a pre-existing “objective” value independent of the measurement. We extend this notion to GPTs and introduce a quantitative measure of intersubjectivity as follows.
Definition 1.
A measurement $A=(a_{x})_{x\in X}$ is called $\alpha$-intersubjective if
$\forall B=(b_{x,x^{\prime}})_{x,x^{\prime}\in X}\in\mathrm{JM}(A,A),\quad\sum_{x\in X}b_{x,x}\geq\alpha 1_{S}.$ (1)
A measurement is said to be intersubjective when it is $1$-intersubjective.
This definition admits a direct operational interpretation: a measurement $A$ is $\alpha$-intersubjective if, whenever two observers perform it simultaneously, they agree on the outcome with a probability of at least $\alpha$.
To analyze intersubjectivity, we also introduce a closely related concept: sharpness. Sharp effects were proposed by Gudder [10, 11, 12] as a generalization of projection operators in quantum theory, motivated by quantum logic [53, 5]. Extending this idea, we define the sharpness of a measurement as follows.
Definition 2.
A measurement $A=(a_{x})_{x\in X}$ is called $\alpha$-sharp if, for any $x\neq x^{\prime}\in X$,
$\forall b\in\mathcal{E}_{S},\quad(b\leq a_{x}\text{ and }b\leq a_{x^{\prime}})\implies b\leq(1-\alpha)1_{S}.$ (2)
A measurement is said to be sharp when it is $1$-sharp.
We call an effect $a$ sharp if the two-outcome measurement $A:=(a,1_{S}-a)$ is sharp, thereby recovering Gudder’s original definition [10]. We say that a measurement $A=(a_{x})_{x\in X}$ is elementwise sharp if $a_{x}$ is sharp for all $x\in X$.
We now show that the conditions of intersubjectivity and sharpness, although conceptually distinct at first glance, are, in fact, equivalent in any physical theory:
Proposition 1.
A measurement is intersubjective if and only if it is sharp.
Moreover, this equivalence admits a quantitative refinement, which is presented in the End Matter.
Characterization of PVMs.—
Although Ozawa [23] proved that all PVMs are intersubjective and that not all POVMs are intersubjective, it remained an open question whether there exists an intersubjective POVM other than PVMs. By applying Proposition 1 to quantum theory, we find the following theorem.
Theorem 2.
A POVM $A=(a_{x})_{x\in X}$ is intersubjective if and only if, for any $x\neq x^{\prime}\in X$,
$\mathrm{supp}(a_{x})\cap\mathrm{supp}(a_{x^{\prime}})=\{0\}.$ (3)
[…]
To fill this gap, we introduce the notion of completely intersubjective measurements by focusing on the coarse-graining of measurements. A measurement $B=(b_{y})_{y\in Y}$ is called a coarse-graining of a measurement $A=(a_{x})_{x\in X}$ if there exists a map $\pi:X\to Y$ such that $\sum_{x\in\pi^{-1}(\{y\})}a_{x}=b_{y}$ for all $y\in Y$.
Definition 3.
A measurement is called completely ($\alpha$-)intersubjective if every coarse-graining of the measurement is ($\alpha$-)intersubjective.
In quantum theory, an effect is sharp if and only if it is represented by a projection operator. Using this fact, we obtain the following characterization.
Theorem 3.
A POVM is a PVM if and only if it is completely intersubjective.
This theorem establishes that complete intersubjectivity provides an operational characterization of PVMs, which represent physical observables in quantum theory. Complete intersubjectivity requires that measurement outcomes be shared by all observers, independently of the measurement resolution; such observer-independent objectivity is a basic property expected of any “physical observable.” Consequently, in GPTs where PVMs are not defined, completely intersubjective measurements naturally arise as a reasonable candidate for the proper definition of “physical observables.”
Furthermore, the quantification of complete intersubjectivity satisfies properties desirable for a measure of “observable-likeness” of measurements: it admits a direct operational interpretation and completely characterizes PVMs in quantum theory.
Quantitative properties of intersubjectivity.—
Next, we discuss the basic properties of the degree of intersubjectivity. First, we describe its behavior under joint measurements and probabilistic mixtures:
  • (i)
    If measurements $A$ and $B$ are $\alpha$- and $\beta$-intersubjective, respectively, then any joint measurement $C\in\mathrm{JM}(A,B)$ is $(\alpha+\beta-1)$-intersubjective.
  • (ii)
    Consider a probabilistic mixture $C=\lambda A+(1-\lambda)B$ with $\lambda\in(0,1]$, where $A=(a_{x})_{x\in X}$ and $B=(b_{x})_{x\in X}$. If $A$ is not (completely) $\alpha$-intersubjective, then $C$ is not (completely) $(1-\lambda(1-\alpha))$-intersubjective.
Here, the probabilistic mixture of two measurements $A=(a_{x})_{x\in X}$ and $B=(b_{x})_{x\in X}$ is defined by $\lambda A+(1-\lambda)B:=(\lambda a_{x}+(1-\lambda)b_{x})_{x\in X}$ for $\lambda\in[0,1]$. The proof is given in the Supplemental Material footnotemark: .
As a special case of statement (i), a joint measurement of intersubjective measurements must be intersubjective. However, a joint measurement of two completely intersubjective measurements is not always completely intersubjective in general, even though this holds in classical and quantum theories [54, 55, 56]. An explicit counterexample is presented in Example 4 in the End Matter.
Statement (ii) implies that adding noise does not improve the intersubjectivity of a measurement. This suggests that the quantification of (complete) intersubjectivity can also be interpreted as quantitative indicators of “noiselessness” of measurements. Importantly, this interpretation aligns with the historical perspective: PVMs have traditionally been regarded as perfectly accurate, while POVMs are generalized measurements that may incorporate noise. Accordingly, the quantification of “PVM-likeness” has also been discussed in the context of measuring noise [18, 19, 20, 21, 22].
These observations demonstrate the importance of the degree of (complete) intersubjectivity. However, evaluating this quantity for a given measurement is generally nontrivial. We now present results for several simple cases. A coin-tossing measurement $A=(\lambda_{x}1_{S})_{x\in X}$ is (completely) $\alpha$-intersubjective if and only if
$\alpha\leq\max\Bigl\{2\max_{x\in X}\lambda_{x}-1,0\Bigr\}.$ (4)
In a classical system $S$, a measurement $A=(a_{x})_{x\in X}$ is (completely) $\alpha$-intersubjective if and only if
$\alpha\leq\inf_{s\in\mathrm{ext}(S)}\max\Bigl\{2\max_{x\in X}a_{x}(s)-1,0\Bigr\}.$ (5)
In a qubit system, an unbiased two-outcome measurement [19] $A=\bigl(\frac{1}{2}(I+\bm{\lambda}\cdot\bm{\sigma}),\frac{1}{2}(I-\bm{\lambda}\cdot\bm{\sigma})\bigr)$, where $\bm{\sigma}:=(\sigma_{1},\sigma_{2},\sigma_{3})$ denotes the Pauli matrices, is (completely) $\alpha$-intersubjective if and only if
[…]
The proofs of these results are provided in the Supplemental Material footnotemark: .
Relationships among conditions.—
We now explore the relationships among similar conditions for measurements: intersubjectivity, complete intersubjectivity, elementwise sharpness, and extremality in the space of measurements. In classical theory, all of these conditions are equivalent, whereas in GPTs, they can diverge (Fig. 2).
ISExtESCIS
(a) GPTs
ISExtES = CIS
(b) Quantum theory
IS = Ext = ES = CIS
(c) Classical theory
Figure 2: Inclusion relation among the set of intersubjective (IS), completely intersubjective (CIS), and elementwise sharp (ES) measurements and the extreme points of the space of measurements (Ext) in (a) GPTs, (b) quantum theory and (c) classical theory.
From the definitions, it is straightforward that completely intersubjective $\implies$ elementwise sharp $\implies$ intersubjective. Moreover, it follows from the result in [57], which relates extremality to the uniqueness of joint measurements, that every extremal measurement $A\in\mathrm{ext}(\mathcal{M}_{S}(X))$ is intersubjective. This observation is significant: the Krein–Milman theorem implies that any measurement can be realized as a convex mixture of intersubjective measurements, which also indicates that intersubjective measurements can be regarded as “noiseless” or “accurate” measurements, while general measurements may be understood as their error-tolerant generalizations.
In quantum theory, in addition to the above general relations, the implications elementwise sharp $\implies$ completely intersubjective $\implies$ extremal hold; in particular, complete intersubjectivity and elementwise sharpness are both equivalent to being a PVM.
Apart from the inclusion relations stated above, no other relations hold in general. For example, in quantum theory there exist (i) extremal POVMs that are not PVMs, and (ii) intersubjective POVMs that are not extremal. In more general systems, there exist measurements that are (iii) completely intersubjective but not extremal, (iv) elementwise sharp and extremal but not completely intersubjective, and (v) elementwise sharp but neither extremal nor completely intersubjective. Explicit counterexamples are provided in the End Matter.
Characterization of classical theory.—
One of the prominent results from studies of GPTs is that certain features once regarded as specific mysteries of quantum theory, such as the existence of incompatible measurements [28, 29] and entanglement [30, 31], are, in fact, common to all non-classical theories. Building on this perspective, we show that the gap between intersubjectivity and complete intersubjectivity provides a novel operational characterization of non-classical theory.
Theorem 4.
A finite-dimensional system is classical if and only if all intersubjective measurements on it are completely intersubjective.
For the proof, see the End Matter.
This result demonstrates that in any non-classical theory, the objectivity of measurement outcomes can depend on the resolution of the measurement and may be lost through coarse-graining. This phenomenon can be illustrated by the following paradoxical thought experiment. Suppose that there are three boxes, and a particle is located in one of them. Two observers, Alice and Bob, attempt to find where the particle is. Assume that the measurement of the location of the particle is intersubjective but not completely intersubjective. Then the following counterintuitive situation arises: if they open all three boxes and simultaneously observe the location of the particle, the intersubjectivity ensures the consensus of their observation; if they open only one box and check whether the particle is inside that box, they may arrive at different conclusions because the measurement is not completely intersubjective. Example 1 in the End Matter realizes this thought experiment within quantum theory.
Advantage in information processing.—
Beyond the operational insights presented so far, we now clarify the significance of (complete) intersubjectivity in specific information-processing tasks: state tomography [38, 39, 40] and state discrimination [41, 42, 43, 37], which play a fundamental role in (quantum) information processing.
State tomography is the informational task of determining an unknown state from the statistics of measurement outcomes. A set of measurements $\mathcal{M}$ is called tomographically complete if it separates the state space; that is, for each $s_{1}\neq s_{2}\in S$, there exists a measurement $A=(a_{x})_{x\in X}\in\mathcal{M}$ and $x\in X$ such that $a_{x}(s_{1})\neq a_{x}(s_{2})$. If $\mathcal{M}$ is tomographically complete, state tomography can be performed using measurements solely from $\mathcal{M}$.
State discrimination is the task of distinguishing a family of states through a single-shot measurement: given a finite ensemble of states $(s_{x},p_{x})_{x\in X}$, where state $s_{x}$ is prepared with probability $p_{x}$, the goal is to identify which state has been prepared. One performs a measurement $A=(a_{x})_{x\in X}$ on the given state and guesses that the prepared state was $s_{x}$ when outcome $x$ is obtained. The success probability of this task is given by $\sum_{x\in X}p_{x}a_{x}(s_{x})$. A family of states $(s_{x})_{x\in X}$ is said to be perfectly distinguished by a measurement $A=(a_{x})_{x\in X}$ if $a_{x}(s_{x})=1$ holds for all $x\in X$.
We observe the following:
  • (i)
    The set of all completely intersubjective measurements is tomographycally complete.
  • (ii)
    For any given ensemble of states, there always exists an intersubjective measurement that is optimal, i.e., that maximizes the success probability of the discrimination task.
  • (iii)
    Every completely intersubjective measurement admits a family of states that is perfectly distinguished by it.
Statement (i) follows from the fact that every extremal two-outcome measurement is sharp [12], while (ii) follows from Bauer’s maximum principle and the implication extremal $\implies$ intersubjective. Statement (iii) follows from the implication completely intersubjective $\implies$ elementwise sharp, along with the fact that for any sharp effect $a\neq 0_{S}$, there exists a state $s$ such that $a(s)=1$ [12]. We note, however, that unlike in the quantum case, a perfectly distinguishable family of states need not be perfectly distinguished by a completely intersubjective measurement, as demonstrated by Example 4 in the End Matter.
When defining a concept in GPTs, one must keep in mind that instances of it may arise only rarely in generic systems [16, 13]. Despite this concern, the above results demonstrate the abundant existence of (completely) intersubjective measurements.
Conclusion.—
In this work, we have investigated the operational foundations of measurements corresponding to physical observables in quantum and generalized probabilistic theories. We have extended Ozawa’s notion of intersubjectivity and introduced complete intersubjectivity, which requires that any coarse-graining of a measurement preserves intersubjectivity. In quantum theory, this property completely characterizes projective measurements (PVMs), providing a natural operational criterion for a measurement to represent a physical observable.
We have quantified intersubjectivity and complete intersubjectivity, providing operational measures of “observable-likeness” and “noiselessness” of measurements. Our results also reveal that the gap between intersubjectivity and complete intersubjectivity serves as a distinguishing feature of non-classical theories, providing a new operational criterion to characterize classical systems.
Finally, we have shown that sufficiently many (completely) intersubjective measurements exist for state tomography and state discrimination tasks in any GPT. This highlights their conceptual and practical roles: not only do they ensure shared outcomes among observers, but they also provide resources for information processing, even beyond quantum theory. These results demonstrate that intersubjectivity unifies foundational and operational perspectives on measurements.
Future work may explore several directions. First, an intriguing question is whether quantum theory can be characterized by additional inclusion relations among the notions related to intersubjectivity. Second, it is also natural to ask whether intersubjectivity admits a resource-theoretic formulation and how it relates to measurement incompatibility. Finally, further developing an operational understanding of broad topics in traditional observable-based physics may not only deepen its conceptual foundations but also provide a basis for formulating and exploring diverse physical theories within the framework of GPTs.
Acknowledgments.—
The authors thank the Yukawa Institute for Theoretical Physics at Kyoto University, where this work was initiated during the YITP-W-25-11. We thank Hiroatsu Okamitsu, Masanao Ozawa, Ryo Takakura, Kaito Watanabe, and Ludovico Lami for fruitful discussions. S. U. was supported by Forefront Physics and Mathematics Program to Drive Transformation, a World-leading Innovative Graduate Study Program, the University of Tokyo. K. O. was supported by Advanced Basic Science Course, a World-leading Innovative Graduate Study Program, the University of Tokyo. H. A. was supported by JSPS KAKENHI Grant Number 25KJ0043 and 26K17031.
Author contributions.—
S. U. and K. O. contributed equally to this work.
Supplemental Material
Appendix S1 Characterization of intersubjective POVMs in infinite-dimensional quantum systems
We present a characterization of intersubjective POVMs in quantum systems represented by Hilbert spaces that are not necessarily finite-dimensional.
Theorem S.1 (Infinite-dimentional version of Theorem 2).
A POVM $A=(a_{x})_{x\in X}$ is intersubjective if and only if, for any $x\neq x^{\prime}\in X$,
$\mathrm{supp}(a_{x})\cap\mathrm{supp}(a_{x^{\prime}})=\{0\}.$ (S1.8)
Proof.
By Proposition 1, it suffices to show that for any positive operators $a,b\geq 0$, the following two conditions are equivalent:
  • (i)
    $\mathrm{supp}(a)\cap\mathrm{supp}(b)=\{0\}.$
  • (ii)
    For any positive operator $c\geq 0$, if $c\leq a$ and $c\leq b$, then $c=0$.
First, assume (i). Let $c\geq 0$ satisfy $c\leq a$ and $c\leq b$. Then $\mathrm{supp}(c)\subseteq\mathrm{supp}(a)$ and $\mathrm{supp}(c)\subseteq\mathrm{supp}(b)$, and hence, $\mathrm{supp}(c)={0}$, which implies $c=0$.
Next, assume (ii), and suppose, to the contrary, that $\mathrm{supp}(a)\cap\mathrm{supp}(b)\neq\{0\}.$ Let
$a=\int_{0}^{\infty}\lambda\,de_{a}(\lambda),\quad b=\int_{0}^{\infty}\lambda\,de_{b}(\lambda)$
be the spectral decompositions of $a$ and $b$, respectively. For $\varepsilon>0$, define
$e_{a,\varepsilon}:=\int_{\varepsilon}^{\infty}de_{a}(\lambda),\quad e_{b,\varepsilon}:=\int_{\varepsilon}^{\infty}de_{b}(\lambda).$
As $\varepsilon\to 0^{+}$, the projections $e_{a,\varepsilon}$ and $e_{b,\varepsilon}$ converge in the strong operator topology to the support projections $p_{\mathrm{supp}(a)}$ and $p_{\mathrm{supp}(b)}$. Consequently,
$e_{a,\varepsilon}\wedge e_{b,\varepsilon}\xrightarrow{\varepsilon\to 0^{+}}p_{\mathrm{supp}(a)}\wedge p_{\mathrm{supp}(b)}>0,$
where $p\wedge q$ denotes the greatest lower bound of $p$ and $q$ in the complete lattice of projections on the Hilbert space. Hence, there exists $\varepsilon>0$ such that $e_{a,\varepsilon}\wedge e_{b,\varepsilon}>0$. Fix such an $\varepsilon$ and define $c:=\varepsilon(e_{a,\varepsilon}\wedge e_{b,\varepsilon})$. Then $c\leq\varepsilon e_{a,\varepsilon}\leq a$ and $c\leq\varepsilon e_{b,\varepsilon}\leq b$, while $c>0$. This contradicts (ii). ∎
Remark.
The above argument extends straightforwardly to more general quantum theories in which a system is described by a von Neumann algebra.
Appendix S2 Proofs of quantitative properties of intersubjectivity
In this appendix, we prove the quantitative properties of intersubjectivity stated in the main text. We begin with its behavior under joint measurements and probabilistic mixtures.
Proposition S.2.
  • (i)
    Suppose that two measurements $A$ and $B$ are $\alpha$- and $\beta$-intersubjective, respectively. Then any joint measurement $C\in\mathrm{JM}(A,B)$ is $(\alpha+\beta-1)$-intersubjective.
  • (ii)
    Let $C=\lambda A+(1-\lambda)B$ with $\lambda\in(0,1]$ be a probabilistic mixture of two measurements $A=(a_{x})_{x\in X}$ and $B=(b_{x})_{x\in X}$. If $A$ is not (completely) $\alpha$-intersubjective, then $C$ is not (completely) $(1-\lambda(1-\alpha))$-intersubjective.
Proof.
  • (i)
    Let $D=(d_{x,y,x^{\prime},y^{\prime}})_{(x,y),(x^{\prime},y^{\prime})\in X\times Y}\in\mathrm{JM}(C,C)$. Define two measurements $E=(e_{x,x^{\prime}})_{x,x^{\prime}\in X}$ and $F=(f_{y,y^{\prime}})_{y,y^{\prime}\in Y}$ by
    $e_{x,x^{\prime}}:=\sum_{y,y^{\prime}}d_{x,y,x^{\prime},y^{\prime}},\quad f_{y,y^{\prime}}:=\sum_{x,x^{\prime}}d_{x,y,x^{\prime},y^{\prime}}.$
    so that $E\in\mathrm{JM}(A,A)$ and $F\in\mathrm{JM}(B,B)$. The intersubjectivity of $A$ and $B$ gives
    $\sum_{x\in X}e_{x,x}\geq\alpha 1_{S},\quad\sum_{y\in Y}f_{y,y}\geq\beta 1_{S}.$
    Therefore,
    $\sum_{(x,y)\in X\times Y}d_{x,y,x,y}$ $=\sum_{x,y,y^{\prime}}d_{x,y,x,y^{\prime}}+\sum_{x,y,x^{\prime}}d_{x,y,x^{\prime},y}$
    $\qquad{}-\sum_{\begin{subarray}{c}x,y,x^{\prime},y^{\prime}\\ x=x^{\prime}\text{ or }y=y^{\prime}\end{subarray}}d_{x,y,x^{\prime},y^{\prime}}$
    $\geq(\alpha+\beta-1)1_{S}.$
    This completes the proof.
  • (ii)
    We first prove the statement for intersubjectivity. Suppose that $C$ is $(1-\lambda(1-\alpha))$-intersubjective. Let $D=(d_{x,x^{\prime}})_{x,x^{\prime}\in X}\in\mathrm{JM}(A,A)$ and $E=(e_{x,x^{\prime}})_{x,x^{\prime}\in X}\in\mathrm{JM}(B,B)$. Then we have $\lambda D+(1-\lambda)E\in\mathrm{JM}(C,C)$. Thus,
    $\sum_{x\in X}\bigl(\lambda d_{x,x}+(1-\lambda)e_{x,x}\bigr)\geq(1-\lambda(1-\alpha))1_{S}.$
    Therefore,
    $\sum_{x\in X}d_{x,x}\geq\frac{1-\lambda(1-\alpha)}{\lambda}1_{S}-\frac{1-\lambda}{\lambda}\sum_{x\in X}e_{x,x}\geq\alpha 1_{S}.$
    The statement for complete intersubjectivity follows by applying the same argument to each coarse-graining. ∎
Next, we present detailed calculations of the degree of (complete) intersubjectivity in several simple cases.
Proposition S.3.
A coin-tossing measurement $A=(\lambda_{x}1_{S})_{x\in X}$ is (completely) $\alpha$-intersubjective if and only if
$\alpha\leq\max\Bigl\{2\max_{x\in X}\lambda_{x}-1,0\Bigr\}.$ (S2.9)
Proof.
Without loss of generality, let $X=\{1,\dots,n\}$ and assume that $\lambda_{1}\geq\dots\geq\lambda_{n}$. We distinguish two cases.
Case 1: $\lambda_{1}\geq\frac{1}{2}$. For any $B=(b_{i,j})_{i,j=1}^{n}\in\mathrm{JM}(A,A)$, we have
$\sum_{i=1}^{n}b_{i,i}\geq b_{1,1}=\lambda_{1}1_{S}-\sum_{i=2}^{n}b_{1,i}\geq\lambda_{1}1_{S}-\sum_{i=2}^{n}\lambda_{i}1_{S}=(2\lambda_{1}-1)1_{S}.$
Conversely, define the measurement $B=(b_{i,j})_{i,j=1}^{n}$ by
$b_{i,j}:=\begin{cases}(2\lambda_{1}-1)1_{S}&\text{if }i=j=1,\\ \lambda_{j}1_{S}&\text{if }i=1\text{ and }j\geq 2,\\ \lambda_{i}1_{S}&\text{if }i\geq 2\text{ and }j=1,\\ 0_{S}&\text{otherwise}.\end{cases}$
Then, one verifies that $B\in\mathrm{JM}(A,A)$ and $\sum_{i=1}^{n}b_{i,i}=(2\lambda_{1}-1)1_{S}$. Hence, $A$ is $\alpha$-intersubjective if and only if $\alpha\leq 2\lambda_{1}-1$.
Case 2: $\lambda_{1}<\frac{1}{2}$. We show that there exists $B=(b_{i,j})_{i,j=1}^{n}\in\mathrm{JM}(A,A)$ with $\sum_{i=1}^{n}b_{i,i}=0_{S}$ by induction on $n$. For $n=1,2$, the condition $\lambda_{1}<\frac{1}{2}$ cannot occur. Assume $n\geq 3$. Then there exists $\mu>0$ such that $\sum_{i=2}^{n}\min\{\lambda_{i},\mu\}=1-2\lambda_{1}$. Consider the coin-tossing measurement $C:=\bigl(\frac{\min\{\lambda_{i},\mu\}}{1-2\lambda_{1}}1_{S}\bigr)_{i=1}^{n}$. By the definition of $\mu$, its maximal probability does not exceed $\frac{1}{2}$; hence, by the induction hypothesis, there exists $D=(d_{i,j})_{i,j=2}^{n}\in\mathrm{JM}(C,C)$ with $\sum_{i=2}^{n}d_{i,i}=0_{S}$. Define the measurement $B=(b_{i,j})_{i,j=1}^{n}$ by
$b_{i,j}:=\begin{cases}0_{S}&\text{if }i=j=1,\\ (\lambda_{j}-\min\{\lambda_{j},\mu\})1_{S}&\text{if }i=1\text{ and }j\geq 2,\\ (\lambda_{i}-\min\{\lambda_{i},\mu\})1_{S}&\text{if }i\geq 2\text{ and }j=1,\\ (1-2\lambda_{1})d_{i,j}&\text{otherwise}.\end{cases}$
Then, one verifies that $B\in\mathrm{JM}(A,A)$ and $\sum_{i=1}^{n}b_{i,i}=0_{S}$. Hence, $A$ is $\alpha$-intersubjective if and only if $\alpha\leq 0$.
Combining the two cases, we conclude that $A$ is $\alpha$-intersubjective if and only if $\alpha\leq\max\{2\lambda_{1}-1,0\}$. Since $\max_{x\in X}\lambda_{x}$ is non-decreasing under coarse-graining, the same formula holds for complete intersubjectivity. ∎
Proposition S.4.
In a classical system $S$, a measurement $A=(a_{x})_{x\in X}$ is (completely) $\alpha$-intersubjective if and only if
$\alpha\leq\inf_{s\in\mathrm{ext}(S)}\max\Bigl\{2\max_{x\in X}a_{x}(s)-1,0\Bigr\}.$ (S2.10)
Proof.
For each $s\in\mathrm{ext}(S)$, $(a_{x}(s))_{x\in X}$ is a probability distribution. Constructing a joint distribution $(b_{x,x^{\prime}}(s))_{x,x^{\prime}\in X}$ for it exactly as in the proof of Proposition S.3, we obtain that the maximal value of $\sum_{x\in X}b_{x,x}(s)$ equals $\max\{2\max_{x\in X}a_{x}(s)-1,0\}$. Performing this construction for each $s\in\mathrm{ext}(S)$ and affinely extending it to all states in $S$, we obtain a measurement $B=(b_{x,x^{\prime}})_{x,x^{\prime}\in X}$, which belongs to $\mathrm{JM}(A,A)$. Hence, $A$ is $\alpha$-intersubjective if and only if $\alpha\leq\inf_{s\in\mathrm{ext}(S)}\max\{2\max_{x\in X}a_{x}(s)-1,0\}$. Since $\inf_{s\in\mathrm{ext}(S)}\max_{x\in X}a_{x}(s)$ is non-decreasing under coarse-graining, the same formula holds for complete intersubjectivity. ∎
Proposition S.5.
In a qubit system, an unbiased two-outcome measurement $A=\bigl(\frac{1}{2}(I+\bm{\lambda}\cdot\bm{\sigma}),\frac{1}{2}(I-\bm{\lambda}\cdot\bm{\sigma})\bigr)$ is (completely) $\alpha$-intersubjective if and only if
[…]
Proof.
[…]
$\min_{b}$ $1-\norm{b}_{\infty}$
subject to $0\leq b,\quad b\leq\frac{1}{2}(I+\bm{\lambda}\cdot\bm{\sigma}),\quad b\leq\frac{1}{2}(I-\bm{\lambda}\cdot\bm{\sigma})$
[…]
$\min_{b_{0},\bm{b}}$ $|\displaystyle 1-b_{0}-\absolutevalue{\bm{b}}$
[…]
Figure S.1: Geometric illustration of the optimization problem in Proposition S.5.
Appendix S3 Proofs of Example 3 and Example 4
In this appendix, we provide proofs of Example 3 and Example 4 presented in the End Matter.
Proof of Example 3.
Let
$S:=\left\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}\>\middle|\>-x_{3}\leq x_{1}\leq x_{3},\ -x_{3}\leq x_{2}\leq x_{3},\ x_{3}=1\right\},$
and let $A:=\frac{1}{2}B+\frac{1}{2}C$, where $B=(b_{+},b_{-})$ and $C=(c_{+},c_{-})$ are defined by
$b_{\pm}(x_{1},x_{2},x_{3}):=\frac{1}{2}(x_{3}\pm x_{1}),\quad c_{\pm}(x_{1},x_{2},x_{3}):=\frac{1}{2}(x_{3}\pm x_{2}).$
We show that $A$ is completely intersubjective but not extremal. It is clear that $A$ is not extremal. Hence, it suffices to show that $A$ is sharp. Suppose that an effect $d$ satisfies
$d\leq\frac{1}{2}b_{+}+\frac{1}{2}c_{+},\quad d\leq\frac{1}{2}b_{-}+\frac{1}{2}c_{-}.$
Evaluating these inequalities at the states (1,1,1)(-1,-1,1) and $(1,1,1)$, respectively, gives d(1,1,1)=d(1,1,1)=0d(-1,-1,1)=d(1,1,1)=0. Moreover, since $d$ is affine, we have d(1,1,1)+d(1,1,1)=d(1,1,1)+d(1,1,1)=0d(1,-1,1)+d(-1,1,1)=d(-1,-1,1)+d(1,1,1)=0. By positivity of $d$, this implies d(1,1,1)=d(1,1,1)=0d(1,-1,1)=d(-1,1,1)=0. Therefore, $d=0_{S}$, and hence $A$ is sharp. ∎
Proof of Example 4.
Let
$S:=\left\{(x_{1},x_{2},x_{3},x_{4},x_{5})\in\mathbb{R}^{5}\>\middle|\>0\leq x_{i}\ (i=1,\dots,5),\ x_{5}\leq x_{1}+x_{4},\ x_{5}\leq x_{2}+x_{3},\ x_{1}+x_{2}+x_{3}+x_{4}=1\right\}.$
Let $A=(a_{1},a_{2},a_{3},a_{4})$, where
$a_{i}(x_{1},x_{2},x_{3},x_{4},x_{5}):=x_{i}\quad(i=1,2,3,4).$
We also define $B:=(a_{1}+a_{2},a_{3}+a_{4})$, $C:=(a_{1}+a_{3},a_{2}+a_{4})$, and $s_{i}:=(\delta_{1,i},\delta_{2,i},\delta_{3,i},\delta_{4,i},0)\in S$ for $i=1,2,3,4$.
We first show that $A$ is elementwise sharp and extremal, but not completely intersubjective. To prove elementwise sharpness, suppose that an effect $e$ satisfies $e\leq a_{1}$ and $e\leq 1_{S}-a_{1}$. Then $e$ must be of the form
$e(x_{1},x_{2},x_{3},x_{4},x_{5})=\lambda x_{5}$
for some constant $\lambda$. Positivity of $e$ gives $\lambda\geq 0$, while evaluating $e\leq a_{1}$ at $\bigl(0,0,\frac{1}{2},\frac{1}{2},\frac{1}{2}\bigr)\in S$ gives $\lambda\leq 0$. Thus $\lambda=0$, and hence $e=0_{S}$. Therefore $a_{1}$ is a sharp effect. The same argument applies to $a_{2}$, $a_{3}$, and $a_{4}$.
The extremality of $A$ follows from the fact that $a_{1},a_{2},a_{3},a_{4}$ are linearly independent indecomposable effects.
On the other hand, $A$ is not completely intersubjective. Indeed, the effect $e$, defined by
$e(x_{1},x_{2},x_{3},x_{4},x_{5}):=x_{5},$
satisfies $e\leq a_{1}+a_{4}$ and $e\leq a_{2}+a_{3}$. Thus the coarse-graining $D:=(a_{1}+a_{4},a_{2}+a_{3})$ of $A$ is not sharp, and hence $A$ is not completely intersubjective.
We next show that $B$ and $C$ are completely intersubjective. Suppose that an effect $e$ satisfies $e\leq a_{1}+a_{2}$ and $e\leq a_{3}+a_{4}$. Then $e$ must be of the form
$e(x_{1},x_{2},x_{3},x_{4},x_{5})=\lambda x_{5}$
for some constant $\lambda$. Positivity of $e$ gives $\lambda\geq 0$, while evaluating $e\leq a_{1}+a_{2}$ at $\bigl(0,0,\frac{1}{2},\frac{1}{2},\frac{1}{2}\bigr)\in S$ gives $\lambda\leq 0$. Thus $\lambda=0$, and hence $e=0_{S}$. Therefore, $B$ is sharp, and hence completely intersubjective. The same argument applies to $C$.
Finally, we show that $(s_{1},s_{2},s_{3},s_{4})$ is perfectly distinguished only by $A$. Let $A^{\prime}=(a^{\prime}_{1},a^{\prime}_{2},a^{\prime}_{3},a^{\prime}_{4})$ be a measurement that perfectly distinguishes these states. Then each $a^{\prime}_{i}$ must be of the form
$a^{\prime}_{i}(x_{1},x_{2},x_{3},x_{4},x_{5})=x_{i}+\lambda_{i}x_{5}$
for some constant $\lambda_{i}$. Positivity of $a^{\prime}_{i}$ implies $\lambda_{i}\geq 0$ for each $i$. Since $A^{\prime}$ is a measurement, we have $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}=0$. Therefore $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=0$, which implies $A^{\prime}=A$. ∎
Appendix S4 Results for continuous-outcome measurements
Several of our results presented in the main text remain valid for continuous-outcome measurements. In this appendix, we extend our analysis to this more general setting.
A continuous-outcome measurement is described by an Effect-Valued Measure (EVM) defined on a measurable space $(X,\Sigma_{X})$. Here, $X$ denotes the outcome space and $\Sigma_{X}$ the $\sigma$-algebra of its measurable subsets. An EVM on $X$ is a map $A:\Sigma_{X}\to\mathcal{E}_{S}$ such that [2]:
  1. 1.
    $A(X)=1_{S}$;
  2. 2.
    For any countable collection of mutually disjoint sets $\{U_{n}\}_{n\in\mathbb{N}}\subseteq\Sigma_{X}$,
    $A\Biggl(\bigcup_{n\in\mathbb{N}}U_{n}\Biggr)=\sum_{n\in\mathbb{N}}A(U_{n}),$ (S4.12)
    i.e., $A\bigl(\bigcup_{n\in\mathbb{N}}U_{n}\bigr)(s)=\sum_{n\in\mathbb{N}}A(U_{n})(s)$ for every state $s\in S$.
For a state $s\in S$, the map $U\mapsto A(U)(s)$ defines a probability measure on $X$, representing the outcome statistics of the measurement. Finite-outcome measurements, which we discussed in the main text, correspond to the special case in which $X$ is a finite set equipped with the discrete $\sigma$-algebra. As in the finite-outcome case, we denote by $\mathcal{M}_{S}(X)$ the set of all EVMs on $X$.
A measurement $C\in\mathcal{M}_{S}(X\times Y)$ is called a joint measurement of a pair of measurements $A\in\mathcal{M}_{S}(X)$ and $B\in\mathcal{M}_{S}(Y)$ if its marginals reproduce $A$ and $B$, that is,
$\forall U\in\Sigma_{X},\quad C(U\times Y)=A(U),\quad\text{and}\quad\forall V\in\Sigma_{Y},\quad C(X\times V)=B(V).$ (S4.13)
We denote the set of all joint measurements of $A$ and $B$ by $\mathrm{JM}(A,B)$.
A measurement $B\in\mathcal{M}_{S}(Y)$ is called a coarse-graining of a measurement $A\in\mathcal{M}_{S}(X)$ if there exists a measurable map $\pi:X\to Y$ such that
$\forall V\in\Sigma_{Y},\quad A\bigl(\pi^{-1}(V)\bigr)=B(V).$ (S4.14)
We now introduce the notion of (complete) intersubjectivity for continuous-outcome measurements.
Definition S.1.
A measurement $A\in\mathcal{M}_{S}(X)$ is said to be intersubjective if $\mathrm{JM}(A,A)$ is a singleton, that is, if every $B\in\mathrm{JM}(A,A)$ coincides with the canonical joint measurement:
$\forall U_{1},U_{2}\in\Sigma_{X},\quad B(U_{1}\times U_{2})=A(U_{1}\cap U_{2}).$ (S4.15)
We say that $A$ is completely intersubjective if every coarse-graining of $A$ is intersubjective.
Lemma S.6.
A measurement $A\in\mathcal{M}_{S}(X)$ is completely intersubjective if and only if $A(U)$ is a sharp effect for every $U\in\Sigma_{X}$.
Proof.
If $A$ is completely intersubjective, then for any $U\in\Sigma_{X}$, the two-outcome coarse-graining $B:=(A(U),A(X\setminus U))$ is intersubjective. Therefore, $A(U)$ is a sharp effect.
Conversely, suppose that $A(U)$ is a sharp effect for every $U\in\Sigma_{X}$. Let $B\in\mathcal{M}_{S}(Y)$ be a coarse-graining of $A$ via a measurable map $\pi:X\to Y$, and let $C\in\mathrm{JM}(B,B)$. For any $V_{1},V_{2}\in\Sigma_{Y}$, we have
$C(V_{1}\times V_{2})\leq B(V_{1})=A\bigl(\pi^{-1}(V_{1})\bigr),\quad C(V_{1}\times V_{2})\leq B(V_{2})=A\bigl(\pi^{-1}(V_{2})\bigr).$
In particular, if $V_{1}\cap V_{2}=\emptyset$, then $\pi^{-1}(V_{1})\cap\pi^{-1}(V_{2})=\emptyset$, and the sharpness of $A\bigl(\pi^{-1}(V_{1})\bigr)$ implies $C(V_{1}\times V_{2})=0_{S}$. For arbitrary $V_{1},V_{2}\in\Sigma_{Y}$, it then follows that
$C(V_{1}\times V_{2})$ $=C((V_{1}\cap V_{2})\times V_{2})+C((V_{1}\setminus V_{2})\times V_{2})$
$=C((V_{1}\cap V_{2})\times V_{2})$
$\leq B(V_{1}\cap V_{2}).$
Summing this inequality over all choices of complements of $V_{1}$ and $V_{2}$, both sides add up to $1_{S}$. Hence, equality must hold term by term:
$\forall V_{1},V_{2}\in\Sigma_{Y},\quad C(V_{1}\times V_{2})=B(V_{1}\cap V_{2}).$
Since the sets of the form $V_{1}\times V_{2}$ generate the $\sigma$-algebra on $Y\times Y$, Carathéodory’s extension theorem ensures that $C$ coincides with the canonical joint measurement of $B$ with itself. Therefore, $B$ is intersubjective. ∎
From the fact that in quantum theory, an effect is sharp if and only if it is represented by a projection, we obtain the following characterization of PVMs.
Theorem S.7 (Continuous-outcome version of Theorem 3).
A (continuous-outcome) POVM is a PVM if and only if it is completely intersubjective.
Next, we summarize some basic properties of (completely) intersubjective measurements.
Proposition S.8.
Suppose that two measurements $A$ and $B$ are intersubjective. Then any joint measurement $C\in\mathrm{JM}(A,B)$ is also intersubjective.
Proof.
Suppose that $A\in\mathcal{M}_{S}(X)$ and $B\in\mathcal{M}_{S}(Y)$. Let $D\in\mathrm{JM}(C,C)$. Define two measurements $E\in\mathcal{M}_{S}(X\times X)$ and $F\in\mathcal{M}_{S}(Y\times Y)$ by
$E(U_{1}\times U_{2}):=D(U_{1}\times Y\times U_{2}\times Y),\quad F(V_{1}\times V_{2}):=D(X\times V_{1}\times X\times V_{2}),$
so that $E\in\mathrm{JM}(A,A)$ and $F\in\mathrm{JM}(B,B)$. The intersubjectivity of $A$ and $B$ gives
$\forall U_{1},U_{2}\in\Sigma_{X},\quad E(U_{1}\times U_{2})=A(U_{1}\cap U_{2}),\quad\text{and}\quad\forall V_{1},V_{2}\in\Sigma_{Y},\quad F(V_{1}\times V_{2})=B(V_{1}\cap V_{2}).$
It then follows that
$D(U_{1}\times V_{1}\times U_{2}\times V_{2})$ $\leq D((U_{1}\cap U_{2})\times(V_{1}\cap V_{2})\times U_{2}\times V_{2})+E((U_{1}\setminus U_{2})\times U_{2})+F((V_{1}\setminus V_{2})\times V_{2})$
$=D((U_{1}\cap U_{2})\times(V_{1}\cap V_{2})\times U_{2}\times V_{2})$
$\leq C((U_{1}\cap U_{2})\times(V_{1}\cap V_{2}))$
$=C((U_{1}\times V_{1})\cap(U_{2}\times V_{2})).$
Summing this inequality over all choices of complements of $U_{1}$, $V_{1}$, $U_{2}$, and $V_{2}$, both sides add up to $1_{S}$. Hence, equality must hold term by term:
$\forall U_{1},U_{2}\in\Sigma_{X},\ \forall V_{1},V_{2}\in\Sigma_{Y},\quad D(U_{1}\times V_{1}\times U_{2}\times V_{2})=C((U_{1}\times V_{1})\cap(U_{2}\times V_{2})).$
Since the sets of the form $U_{1}\times V_{1}\times U_{2}\times V_{2}$ generate the $\sigma$-algebra on $X\times Y\times X\times Y$, Carathéodory’s extension theorem ensures that $D$ coincides with the canonical joint measurement of $C$ with itself. Therefore, $C$ is intersubjective. ∎
Proposition S.9.
Let $C=\lambda A+(1-\lambda)B$ with $\lambda\in(0,1]$ be a probabilistic mixture of two measurements $A$ and $B$. If $C$ is (completely) intersubjective, then $A$ is also (completely) intersubjective.
Proof.
We first prove the statement for intersubjectivity. Let $D_{1},D_{2}\in\mathrm{JM}(A,A)$ and $E\in\mathrm{JM}(B,B)$. Then $\lambda D_{1}+(1-\lambda)E$ and $\lambda D_{2}+(1-\lambda)E$ both belong to $\mathrm{JM}(C,C)$. Since $C$ is intersubjective, we must have $\lambda D_{1}+(1-\lambda)E=\lambda D_{2}+(1-\lambda)E$, which implies $D_{1}=D_{2}$. Hence, $A$ is intersubjective.
The statement for complete intersubjectivity follows by applying the same argument to each coarse-graining. ∎
To connect the intersubjectivity of measurements with the extremality, we first establish a useful characterization of extremal measurements in terms of the uniqueness of joint measurements.
Lemma S.10.
For a measurement $A\in\mathcal{M}_{S}(X)$, the following three conditions are equivalent:
  • (i)
    AA is an extreme point of $\mathcal{M}_{S}(X)$.
  • (ii)
    For every measurement $B$, if a joint measurement of $A$ and $B$ exists, it is unique.
  • (iii)
    There exists a nontrivial coin-tossing measurement $B$ such that the joint measurement of $A$ and $B$ is unique. Here, a measurement $B\in\mathcal{M}_{S}(Y)$ is called a coin-tossing measurement if there exists a probability measure $\mu$ on $Y$ such that $B(V)=\mu(V)1_{S}$ for all $V\in\Sigma_{Y}$. It is called nontrivial if there exists $V\in\Sigma_{Y}$ with $0<\mu(V)<1$.
Proof.
[…]
$D_{1}(U)$ $=C_{1}(U\times V_{0})+C_{2}(U\times(Y\setminus V_{0}))=A(U)+C_{1}(U\times V_{0})-C_{2}(U\times V_{0}),$
$D_{2}(U)$ $=C_{1}(U\times(Y\setminus V_{0}))+C_{2}(U\times V_{0})=A(U)-C_{1}(U\times V_{0})+C_{2}(U\times V_{0}).$
Then $A=\frac{1}{2}D_{1}+\frac{1}{2}D_{2}$. Since $A$ is extremal in $\mathcal{M}_{S}(X)$, we must have $D_{1}=D_{2}$, implying that $C_{1}(U\times V_{0})=C_{2}(U\times V_{0})$ for all $U\in\Sigma_{X}$. As this holds for arbitrary $V_{0}\in\Sigma_{Y}$, we conclude $C_{1}=C_{2}$. Hence, the joint measurement is unique.
[…]
$D(U\times V):=\mu(V\cap V_{0})C_{1}(U)+\mu(V\setminus V_{0})C_{2}(U).$
Then $D$ is a joint measurement of $A$ and $B$. By the uniqueness assumption, this joint measurement must coincide with the canonical one: $D(U\times V)=A(U)B(V)=\mu(V)A(U)$. In particular, evaluating $D$ on $U\times V_{0}$, we conclude that $C_{1}(U)=A(U)$ for all $U\in\Sigma_{X}$. Hence, $A$ is an extreme point of $\mathcal{M}_{S}(X)$. ∎
Remark.
Lemma S.10 extends the result of Ref. [57] to continuous-outcome measurements. The same argument also applies when $A$ is taken to be a general channel rather than a measurement. Moreover, condition (ii) remains valid when $B$ is replaced by an arbitrary channel. In this setting, Lemma S.10 further shows that if $B$ is an extremal channel, then the joint channel, when it exists, is extremal as well. These results provide a characterization of extremal channels and generalize the results of Ref. [56] to the framework of GPTs.
Applying condition (ii) of Lemma S.10 to the case $B:=A$ yields the following proposition:
Proposition S.11.
Every extreme point of $\mathcal{M}_{S}(X)$ is an intersubjective measurement.